40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 321 of 335  Not logged in ELOG logo
ID Date Author Type Category Subject
  733   Thu Jul 24 08:09:26 2008 YoichiUpdateLSCArm cavity g-factor measurement

Quote:

A-ha! Do you always expect the 30Hz signal, don't you?
Because this is the PDH technique.


Yes you are right. I realized this when I was thinking about it in the bed Smile
The 30Hz signal should always be present because the carrier is phase shifted at 30Hz by the cavity length change.
I think the change in the signal ratio between I and Q happened because as the 166MHz sidebands get phase change when they move around the MC transmission peak due to the cavity pole of the MC. It changes the optimal demodulation phase for the 166MHz PDH signal at the AS port.


Quote:

We should be able to see 166MHz sideband resonances using the double demodulated photodetectors. With these, the 33MHz sidebands will be acting as LO when the 166MHz sideband (or mode) resonates. Some modeling may be necessary to determine if the SNR will be good enough to make this worthwhile, however.


I will try, but at 100kHz away from the MC FSR (the number predicted by John's calculation), the transmission of the 166MHz sidebands is very weak. I did not see any signal when I swept it +/- 500kHz. Unfortunately, the Marconi's output level is almost at its maximum. So we don't have much room for increasing the sideband power.
  732   Thu Jul 24 03:08:20 2008 robUpdateLocking+f2 DRMI+2ARMS

rob, john, yoichi

Tonight we tried to move the 166MHz (f2) sideband frequency by changing the settings on the Marconi. Reducing the frequency by 4kHz reduced the amplitude of the 166MHz sidebands, but we were still able to lock the DRMI with the +-f2 sidebands by electronically compensating for the gain decrease, and also to lock the DRMI+2ARMs while resonating the -f2 sideband. No luck with the +f2.

Then we larkily tried increasing the frequency by 4kHz, which ~doubled the f2 sideband transmission through the MC. This means our frequencies/MC length have been mismatched for months. Apparently I explained the level of the f2 sidebands by just imagining that I'd (or someone) had set the modulation depth at that level some time in the past.

It's a miracle any locking worked at all in this state. Once this was done and we worked out a few kinks in the script, adjusting some gains to compensate, we managed to get the DRMI+2ARMS to lock a couple of times while resonating the +f2 sideband. It takes a while, but at least it happens. Tomorrow we'll measure the length of the mode cleaner properly and then try again. No need to vent just yet.
  731   Thu Jul 24 02:57:26 2008 robUpdateLSCArm cavity g-factor measurement

Quote:

So, now I feel that the method for the TEM01 quest should be reconsidered.

If we have any unbalanced resonance for the phase modulation sidebands, the offset of the error signal is to be observed even with the carrier exactly at the resonance. We don't need to shake or move the cavity mirrors.

Presence of the MC makes the things more complicated. Changing the frequency of the modulation that should go throgh the MC is a bit tricky as the detuning produces FM-AM conversion. i.e. The beam incident on the arm cavity may be not only phase modulated but also amplitude modulated. This makes the measurement of the offset described above difficult.

The setup of the abs length measurement (FSR measurement) will be easily used for the measurement of the transverse mode spacings. But it needs some more time to be realized.


We should be able to see 166MHz sideband resonances using the double demodulated photodetectors. With these, the 33MHz sidebands will be acting as LO when the 166MHz sideband (or mode) resonates. Some modeling may be necessary to determine if the SNR will be good enough to make this worthwhile, however.
  730   Thu Jul 24 01:27:00 2008 KojiUpdateLSCArm cavity g-factor measurement

Quote:
I locked the Y-arm and mis-aligned the end mirror in Yaw until the transmission power gets half.
Then I injected a 30Hz sinusoid into the error point of the Y-arm servo to shake the ETMY.
I observed AS166_I and AS166_Q as I changed the 166MHz frequency.


A-ha! Do you always expect the 30Hz signal, don't you?
Because this is the PDH technique.

---------------
Recipe:
You have a carrier and phase modulation sidebands at 166MHz this time.
Inject them into a cavity. Detect the reflection by a photo detector.
Demodulate the photocurrent at 166MHz.

This is the PDH technique.

A 30Hz sinusoid was injected to the error point of the cavity lock.
This means that the cavity length was fluctuated at 30Hz.

We should see the 30Hz signal at the error signal of the 166MHz demodulation, regardless of the tuning of the modulation frequency!
In other words, the 30Hz signal in the demod signal at the 166MHz is also understandable as the beating between the 30Hz sidebands and the 166MHz sidebands.

---------------

So, now I feel that the method for the TEM01 quest should be reconsidered.

If we have any unbalanced resonance for the phase modulation sidebands, the offset of the error signal is to be observed even with the carrier exactly at the resonance. We don't need to shake or move the cavity mirrors.

Presence of the MC makes the things more complicated. Changing the frequency of the modulation that should go throgh the MC is a bit tricky as the detuning produces FM-AM conversion. i.e. The beam incident on the arm cavity may be not only phase modulated but also amplitude modulated. This makes the measurement of the offset described above difficult.

The setup of the abs length measurement (FSR measurement) will be easily used for the measurement of the transverse mode spacings. But it needs some more time to be realized.
  729   Thu Jul 24 01:04:01 2008 robConfigurationLSCIFR2023A (aka MARCONI) settings

Quote:


P.S.: We made a test by changing the frequency of the local oscillator by a little bit and then coming back to the original value. We observed that the phase of the signal can change, so every time this frequency is moved the 3f demod phase need to be retuned.



We discovered this little tidbit in March, and remembered it tonight. Basically we found that whenever you change the frequency on one of these signal generators (and maybe any other setting as well), the phase of the signal can change (it's probably just the sign, but still...), meaning that you when you return settings to their intial value, not everything is exactly as it once was. For most applications, this doesn't matter. For us, where we use one Marconi to demodulate the product of two other Marconis, it means we can easily cause a great deal of grief for ourselves, as the demod phase for the double demod signals can appear to change.

Programmatically, what this means is that every time you touch a Marconi you must elog it. Especially if you change a setting and then put it back.
  728   Wed Jul 23 22:34:07 2008 YoichiUpdateLSCArm cavity g-factor measurement
I tried the same thing as the X-arm to the Y-arm.
I'm puzzled. I found exactly the same behavior as the X-arm in the AS166 demodulated signals, whereas I expected different resonance frequency because of the arm length difference.

Here is more detailed account of the measurement today.

I locked the Y-arm and mis-aligned the end mirror in Yaw until the transmission power gets half.
Then I injected a 30Hz sinusoid into the error point of the Y-arm servo to shake the ETMY.
I observed AS166_I and AS166_Q as I changed the 166MHz frequency.

At 165.977MHz, both AS166_I and AS166_Q showed the 30Hz signal (15cnt p-p).
At 165.981MHz, Only I phase showed the 30Hz signal (40cnt p-p). No signal in Q.
At 165.984MHz, I and Q became the same amplitude again (20cnt p-p).
At 165.987MHz, Only Q phase showed the 30Hz signal (40cnt p-p). No signal in I.

Outside the above range, the signal decreases as the frequency go away. I think this is (at least partly) because the 166MHz sidebands no longer go through the MC at those frequencies.

I then locked the X-arm to the TEM01 mode. I saw exactly the same behavior as described above. This could be the resonance of TEM02 mode. I was expecting to see the resonance of TEM00 mode at the opposite side, but nothing there.

I unlocked the arm cavities and tried the same frequency scan of the 166MHz with one of the end mirrors shaken at 30Hz. I saw no signal at the AS166 port.
I also tried locking Y-arm and shaking the ETMX. No signal.
So it has to be something to do with the cavity resonance.

Since the MC transmission curve for 166MHz is folded in the measurement, it makes the interpretation of the results harder.
  727   Wed Jul 23 21:48:30 2008 robConfigurationGeneralrestore IFO when you're done with it

when you are done with the IFO, please click "Restore last auto-alignment" on the yellow IFO portion of the C1IFO_CONFIGURE.adl screen. Failure to comply will be interpreted as antagonism toward the lock acquisition effort and will be met with excoriation.
  726   Wed Jul 23 18:42:18 2008 JenneUpdatePSLAlignment of AOM
[Rana, Yoichi, Jenne]

Short Version: We are selecting the wrong diffracted beam on the 2nd pass through the AOM (we use the 2nd order rather than the first). This will be fixed tomorrow.

Long Version of AOM activities:

We checked the amount of power going to the AOM, through the AOM on the first pass, and then through the AOM on the second pass, and saw that we get about 50% through on the first pass, but only about 10% on the 2nd pass. Before the AOM=60mW, after the first pass=38mW, after the 2nd pass=4mW. Clearly the alignment through the AOM is really sketchy.

We translated the AOM so the beam goes through the center of the crystal while we align things. We see that we only get the first order beam, which is good. We twiddled the 4 adjust screws on the side of the AOM to maximize the power at the curved mirror for the 1st order of the first pass, which was 49.6mW. We then looked at the DC output of the Reference Cavity's Refl. PD, and saw 150mV on the 'scope. The power measured after the polarizing beam splitter and the next wave plate was still 4mW. Adjusting the curved mirror, we got up to 246mV on the 'scope for the Refl. PD, and 5.16mW after the PBS+Waveplate. We adjusted the 4 side screws of the AOM again, and the tip/tilt of the PBS, and got up to 288mV on the 'scope.

Then we looked at the beam that we keep after the 2nd pass through the AOM, and send to the reference cavity, and we find that we are keeping the SECOND order beam after the second pass. This is bad news. Yoichi and I will fix this in the morning. We checked that we were seeing a higher order beam by modulating the Offset of the MC servo board with a triangle wave, and watching the beam move on the camera. If we were chosing the correct beam, there would be no movement because of the symmetry of 2 passes through the AOM.

I took some sweet video of the beam spot moving, which I'll upload later, if I can figure out how to get the movies off my cell phone.
  725   Wed Jul 23 17:19:48 2008 AlbertoConfigurationComputersMegatron connected
We changed the IP address. Ther new one is 131.215.113.95.

Joe, Alberto


Quote:
Joe, Rana, Alberto,

we found out the password for Megatron so we could log in and set a new one so that now it's the same as that for controls.
The IP address is 131.215.113.59.

We had to switch to another LAN ports to actually connect it.
  724   Wed Jul 23 16:31:02 2008 AlbertoConfigurationComputersMegatron connected
Joe, Rana, Alberto,

we found out the password for Megatron so we could log in and set a new one so that now it's the same as that for controls.
The IP address is 131.215.113.59.

We had to switch to another LAN ports to actually connect it.
  723   Wed Jul 23 13:52:26 2008 SharonUpdate MEDM changes
There is a new MEDM screen now when you go from c1ass>top>pem.
Instead of having 12 "mini filters" screens go to 8 outputs (with the wrong correlation impression from the table), we have a 24X8 matrix.
This matrix is there so you could choose which noise signals you want to send to the adaptive code. When you indicate the number of noise channels you are going to use
on the nAUX option on the screen top, you are choosing the channels 1 to nAUX. Channels 15-22 are the seismic and accelerometers we are now using. (you can see the order in Jenne's post 673).
Hope this will make things clearer.
Attachment 1: matrix
  722   Wed Jul 23 12:42:23 2008 EricSummaryCamerasWeekly Summary
I finally got the ezcaPut command working. The camera code can now talk directly to the EPICS channels. However, after repeated calls of the ezcaPut function, the function begins claim to time out, even though it continues to write values to the channel successfully (EPICS is successfully getting the new value for the channel, but failing to reply back to the program in time, I think). It has seg-faulted once as well, so the stability cannot yet be trusted for running long term. For now, however, it works well enough to test a servo in the short term. The current approach simply uses a terminal running ezcaservo with the pitch and yaw offset channels of the ETMX, as well as the channels that the camera code output to. This hasn't actually been tested since we haven't had enough time with the x-arm locked.

Tested various fixed zoom lens on the camera, since the one we were previously using was too heavy for its mount and likely more expensive than necessary. The 16mm lens gets a good picture of the beam and the optic together, though the beam is a little too small in the picture to reliably fit a gaussian to. The 24mm lens zooms too much to see the whole optic, but the beam profile itself is much clearer. The 24mm lens is currently on the camera.

Scanned the PZT voltage of the PMC across its full offset range to gain a plot of voltage vs intensity. I used DTT's triggered time series response system to measure the outputs of the slow PZT voltage and transmission intensity channels, and used the script triangle wave to drive the PZT ramp channel slowly over its full range (I couldn't get DTT to output to the channel). Clear resonances did appear (PMCScanWide.tif), but the number of data points per peak was far too small reliably fit a lorentzian to (PMCScanSinglePeak.tif). When I decreased the scanning range and increased the time in order to collect a large number of points on a few peaks, the resulting data was too messy to fit to a lorentzian (PMCSlowSinglePeak.tif).
Attachment 1: PMCScanSinglePeak.tif
PMCScanSinglePeak.tif
Attachment 2: PMCScanWide.tif
PMCScanWide.tif
Attachment 3: PMCSlowSinglePeak.tif
PMCSlowSinglePeak.tif
  721   Wed Jul 23 10:49:37 2008 MaxUpdateComputer Scripts / ProgramsWeekly Progress Report
This week I installed the magnetometer. The channels seem to be reading correctly. I'm back to working on noise budget and have added the MICH and will soon add the PRC source. The various source-specific scripts still need to be adjusted and the transfer functions remeasured since they do not match in any reasonable manner the SRD Rana put out in the e-log yesterday.
  720   Wed Jul 23 10:47:05 2008 SharonUpdate Weekly update
This week I spent some time with Alex who updated the adaptive code to save the filter's coeffs all the time, stop when I open the loop, and reload the latest coeffs. when I start it again.
The point was to minimize the adaptation rate. Unfortunately, seems it is making some filters go wild, so it is not in use yet.
After taking some more measurements with the adaptive filter running, I have noticed a new peak in the signal around 22-23 Hz. My first assumption was that this is caused due to internal resonance of MC1 (which is driven when the adaptive code is running, and not when it's not). Therefore, I drove MC1 without the adaptive filter looking for the same peak... which wasn't there.
This sent me back to the adaptive code... Seems there is a matrix in the simulink file of the adaptive filter which doesn't have an MEDM screen. I am now working on making this screen. Once I am done with that, and make sure there is correlation between simulink and MEDM, I'll keep on chasing the peak in the code.
  719   Wed Jul 23 01:42:26 2008 ranaConfigurationPSLFSS RFPD: Examined, "repaired", and re-installed
Rob said that there might be something wrong with the FSS RFPD since the loop gain is so low.
Next time we should just use the Jenne laser on it in-situ and compare with our reference.

We had a 24.5 MHz LSC PD which Rob got from Sam. Sam got it from Rai. I gave it to Rai in Livingston
because it seemed suspicious. Seems fine now. This black box PD had a lower overall response than
the goldbox one we already had. The 2001-2005 era diodes which we got from the Canadian Perkin-Elmer
all had high capacitance and so that's not a surprise.

So the goldbox one was not broken totally.

I found that the offset came from a cracked capacitor. C25 was a yellow thru-hole ceramic 0.1 uF.
Its a surface mount board...don't know why this was like this but there's also no reason it should
have cracked unless it was soldered on with too much heat. I replaced it with a 0.47 uF ceramic
surface mount. Also R24 was a 20 Ohm resistor and L3 was not stuffed?? Removed R24 and put a 1 uH
inductor into L3. This is there so that the input to the MAX4107 is AC coupled.

However, the DC signal that Rob saw was actually because of the cracked C25. It had shorted and was
making a 25 mV signal at the input to the MAX4107 which has a gain of 10. This was producing ~165 mVdc
into a 50 Ohm load and so it could have saturated most mixers. The FSS board, however, has an overly
monstrous level 21 (I think) mixer and so this should not have been an issue. Maybe.

I was able to lock with the 24.5 MHz black box PD but it was not too hard to repair the gold box one
so I did. I tuned it so that the notch is truly at 43 MHz (2x the FSS 21.5 MHz modulation) but because
someone has done this using a hacky cap in parallel with the main PD, I am unable to get the resonant
peak to line up at 21.5 MHz. Its at 23 MHz instead. This loses us ~2 dB in signal. Since the frequency
is so low, we can increase the gain in the MAX4107 by another factor of 3 or so in the future.


So the PD is not our problem. Still worth verifying that the cable is good -- its around 10 miles long!!
And loops around in there with a bunch of other cables. We have an electronic phase shifter so this seems
totally misguided.

The other bad problem is that the mode matching is pretty horrible. Something like 1/3 of the carrier
power doesn't go into the cavity.
FSS TODO:
1) Check cable between RFPD and FSS box for quality. Replace with a good short cable.

2) Using a directional coupler, look at the RFPD output in lock on a scope with 50 Ohm term.
   I suspect its a lot of harmonics because we're overmodulating to compensate for the bad
   mode matching.

3) Purchase translation stages for the FSS mode matching lenses. Same model as the PMC lenses.
   Fix the mode matching.

4) Get the shop to build us up some more bases for the RFPDs on the PSL such as we have for the LSC.
   Right now they're on some cheesy Delrin pedestals. Too soft...

5) Dump the beam reflected off the FSS RFPD with a little piece of black glass or a razor dump.
   Anodized aluminum is no good and wiggles too much.

The attached PDF shows photos of the old and new style PDs. One page 3 there's a wire that I soldered on
as a handle so that we can remove the RF can (occasionally people claim that soldering to the lid screws
up the magnetic shielding magic of the lid. use this as a litmus test of their electronics know-how; its
a tin can - not an orgone box). Pages 4 & 5 are the circuit before I soldered, page 6 the cap after I
tried to remove it, page 7 is the circuit after I put in the new cap, and page 8 is the schematic with
the mark up of the changes.
Attachment 1: Untitled.pdf
Untitled.pdf Untitled.pdf Untitled.pdf Untitled.pdf Untitled.pdf Untitled.pdf Untitled.pdf Untitled.pdf
  718   Tue Jul 22 22:25:31 2008 ranaUpdateLSCLooptickle for existing 40m
John and I have adapted the Stefan-Looptickle model of the 40m upgrade to have the parameters of the old one.
(old one = what we have had for the last 4 years).

Its in the /cvs/cds/caltech/iscmodeling directory on the CDS computers but you can also check it out from the
MIT CVS repo; its part of the whole shebang.

It makes the attached theoretical NB. Feel free to modify it.
Attachment 1: nb.png
nb.png
  717   Tue Jul 22 22:11:58 2008 YoichiUpdateLSCX-arm g factor measurement
Alberto, Yoichi

We measured the g factor of the X-arm by slightly shifting the 166MHz sideband frequency:

We first locked the X-arm to TEM00 mode. Then misaligned the ETMX in yaw a little bit until the transmitted power is a half of the normal value.
In this way, we can expect that TEM01 mode will be resonated in the arm if a sideband with a suitable frequency is introduced.
To add such a sideband, we used the 166MHz EOM. According to John's calculation (ELog entry 690), the TEM01 mode of the 166MHz upper sideband is only about 100kHz away from the resonance. So by changing the 166MHz modulation frequency, we should be able to see the 166MHz upper sideband resonating in the X-arm.
We used the 166MHz PD at the AS to find the resonance.
When we modulated the 166MHz RF frequency by +/- 100kHz, we could see spikes in the AS166_I signal.
Then we fine tuned the RF frequency slowly by hand to find the exact resonant frequency. At that time, the X-arm PDH servo was oscillating at ~480Hz.
So the resonance was determined by maximizing this signal in the AS166_I.
The 166MHz signal was originally at 165.977195 MHz. I found the resonance around 165.985MHz. It is surprisingly close to the original modulation frequency (only 7.3kHz away). This number yields the g factor of 0.626 and the transverse mode interval of 0.285*FSR. I used the arm length of 38.5750m in this calculation. Because of the 480Hz oscillation, it was difficult to precisely determine the resonant frequency. We will try this again tomorrow after mitigating the oscillation.
Although the resonance of the 166MHz upper sideband is located at a lower frequency in John's prediction, we found a resonance at a higher frequency.
This can be interpreted as the discrepancy between the actual g-factor and the designed g-factor.

To confirm what we saw was really an arm cavity resonance, we will try to do the same thing with the arm cavities all mis-aligned.
(We expect no signal in this configuration.)

Appendix: the expected signal from AS166 port when the 166MHz upper sideband passes by a resonance of the arm cavity.
Since the carrier is resonating in the cavity and kept there by the PDH feedback using 33MHz sideband, its phase is virtually fixed at the AS166 port. The lower sideband's phase also does not change much because it is off resonance. The upper sideband get a large phase change when approaching to the resonance. This effectively rotates the modulation angle of the 166MHz sidebands, which was orthogonal to the carrier when off resonance (i.e. phase modulation), to create 166MHz amplitude modulation. Because the sideband axis is rotated, the signal should appear both in I and Q phases.
  716   Tue Jul 22 16:50:09 2008 steveMetaphysicsEnvironmentprofessorial clean up of work bench
Atm1: is showing the spiritual satisfaction after work bench clean up by the professor himself.

Atm2: some items are still waiting to be placed back to their location
Attachment 1: ranaclup.png
ranaclup.png
Attachment 2: cleanup2.png
cleanup2.png
  715   Tue Jul 22 13:16:09 2008 John, RobUpdatePSLFSS open loop transfer function

Quote:
With the common gain slider maxed out the unity gain frequency is 58kHz.

The reference cavity refl diode appears to be okay. RF OUT/ TEST IN transfer function was normal.
There is a ~220mV offset in the RF out. We removed this using a coupler - no change. We also checked the
diode->FSS cable.

Tomorrow I'll take a closer look at the board.


Should note that the UGF of 58kHz was measured with the test cable (from RFPD to board), so the demod phase was presumably sub-optimal.
  714   Tue Jul 22 13:15:14 2008 robUpdatePSLNote from R. Abbott re: the PMC

Quote:
an email from Rich:
Your PZT is broken.

R


Quelle surprise

Frown
  713   Tue Jul 22 11:55:22 2008 ranaUpdatePSLNote from R. Abbott re: the PMC
an email from Rich:
Your PZT is broken.

R
  712   Tue Jul 22 09:24:17 2008 steveUpdatePSLlaser power
Laser power reality of 120 days
Attachment 1: power120d.jpg
power120d.jpg
  711   Tue Jul 22 03:03:22 2008 John, RobUpdatePSLFSS open loop transfer function
With the common gain slider maxed out the unity gain frequency is 58kHz.

The reference cavity refl diode appears to be okay. RF OUT/ TEST IN transfer function was normal.
There is a ~220mV offset in the RF out. We removed this using a coupler - no change. We also checked the
diode->FSS cable.

Tomorrow I'll take a closer look at the board.
  710   Mon Jul 21 19:55:16 2008 rana,jenneConfigurationIOONoise in MC_F
Jenne put the MC board on extender today - its still that way but everything is probably connected (check AO).

We measured the TFs of the DAQ section for MC_F because of how everything looked wrong in the plots Jenne
put in the log earlier. Everything we measured today seemed to jive with the schematic. We also looked up
the original traveler for this board which Betsy filled in years ago: it also is in spec for the DAQ filtering.

So then I looked at the power spectrum of the output signal to the VCO. It had lots of HFC (high frequency crap).
I adjusted the parameters of the FSS (common gain, fast gain, RF phase Astonished) and lowered the MC common gain. This
produced a global minimum in that 4D parameter space.

I think that basically, the FSS gain is too low even with the common gain slider maxed. Having the fast gain up
at 19 dB like I had left it was bad - even though it minimizes the PC control signal, it produces a lot of HFC
up around 100 kHz in MC_F. After John (finally) gets around to measuring the FSS loop we can figure out how to
better tune this. The MC gain then has to be tuned so as to best minimize the HFC given the new FSS gain; there's
basically no coupling from the MC gain to the FSS loop shape so its always best to tune the FSS first. Pleased

The RF phase of the FSS was a mystery - I have no idea why it should do anything and I have never heard of this
and I don't know why I tried it today. But...changing it by ~0.6-0.7 slider units reduced the HFC by another factor
of ~3. Somebody should put this slider into units of degrees.8-)

Here's a table of the changes. Please make these the new nominals:
you asked for:   diff 2008/07/21,13:00 2008/07/22,2:44:16 utc '.*FSS.*|.*MC.*'
LIGO controls: differences, 2008 07/21 13:00:00 utc vs. 2008 07/22 02:44:16 utc
__Epics_Channel_Name______   __Description__________   __value1____     __value2____
C1:IOO-MC_REFL_GAIN                                    22.000000        19.000000
C1:IOO-MC_REFL_OFFSET                                  0.818380         0.818100
C1:PSL-FSS_MGAIN                                       10.000000        30.000000
C1:PSL-FSS_PHCON                                       2.073170         1.413170

The attached plot shows the "SERVO" TNC output of the board; this is supposed to be the same as the voltage going to the
VCO box. So its V/Hz transfer function is flat above 40 Hz. Tomorrow Jenne will post more data and remove the extender
board.

Since I only used an SR785, I only saw noise up to 100 kHz. Its key to use an RF spectrum analyzer when checking out
the FSS and the MC systems.
Attachment 1: SCRN0024.GIF
SCRN0024.GIF
  709   Mon Jul 21 19:48:57 2008 SharonUpdate how tp restart C1ASS
How to restart C1ASS:

1. reboot
2. as root: caltech/target/c1ass:> ./startass
3. no need for root: burtgooey

that's it...
  708   Mon Jul 21 15:52:22 2008 steveBureaucracySAFETYfire alarms test
The fire alarm test and evacuation drill was successfully completed at 13:45 Wednesday, July 16, 2008

Everybody was on time for the 40m meeting.
  707   Mon Jul 21 14:26:11 2008 MaxSummaryPEMAdded Channels
The following channels have been added.

Channel Name DAQ port
C1 : PEM-MAG_BSC_X 27
C1 : PEM-MAG_BSC_Y 28
C1 : PEM-MAG_BSC_Z 29

Jenne and I ran the wires from near the beam splitter chamber (as described in a previous elog) to the rack Y7 and plugged the labeled BNC's into ports 27-29. The computer was c0dcu1. John then restarted the frame builder and Alberto and I restarted the front end of c0dcu1 as per the wiki's instructions. The channels seem to be working. - Max.
  706   Mon Jul 21 11:54:00 2008 JenneUpdateGeneralMC Servo Board
I pulled the MC Servo Board again, to check the components that are on the board, and compare them with the schematics. The filters that I'm interested in on the Fast Path haven't been changed. The high pass filters on the Fast Path have been changed.
Component      Schematic      Actual
---------      ---------      ------
C140           10u            open
C144           10u            open
C149           open           a gray Cap.  value unknown
C141           10u            open
C145           10u            open
R97            1.58K          0
R99            open           1130
R103           open           1130
R100           open           0
R104           100            1130
R98            1.58K          open
R109           367            365

Board is back in, and MC locks.
  705   Mon Jul 21 10:00:42 2008 steveUpdatePEMifo room temp
Five years of temp trend with 5 C fluctuation. The campus chilled water supply was upgraded ~ 1 year ago.

I requested our thermostats to be calibrated.
Attachment 1: temp5y.jpg
temp5y.jpg
  704   Mon Jul 21 09:52:05 2008 SharonUpdate Adaptive code changes
Thanks to Alex, we now save the coefficients of the adaptive filter every cycle. When we choose ENABLE: OFF on the MEDM screen, suppressing the signal to the MC1, we stop and save the last coefficients. When enabling it again, it starts from the last coefficients saved. I will take some measurements today to check how this contributes to the adaptation rate. If you change the number of taps or the number of AUX channels, the coefficients are again set to zero.
  703   Sat Jul 19 19:41:56 2008 YoichiAoGPSLThe author of the entry 702 is Yoichi not Rob
I made a mistake.
  702   Sat Jul 19 19:39:44 2008 robUpdatePSLPMC PZT investigation

Quote:

Quote:
The 64k is the series resistor after the PA85, not the modeled resistance of the PZT itself.

Yes. What I meant was that because the measured voltage across the PZT was a half of the open voltage of the HV amp, the DC impedance of the PZT is expected to be similar to the output impedance of the HV amp. Of course, I don't think the DC impedance of a normal PZT should be such low.
I'm puzzled by the discrepancy between this expected DC impedance and the directly measured impedance by the Fluke DVM (1.5M Ohm).
One possibility is that the PZT leaks current only when a high voltage is applied.
  701   Fri Jul 18 23:24:24 2008 robUpdatePSLPMC PZT investigation

Quote:
I measured the HV coming to the PMC PZT by plugging it off from the PZT and hooking it up to a DVM.
The reading of DVM is pretty much consistent with the reading on EPICS. I got 287V on the DVM when the EPICS says 290V.

Then I used a T to monitor the same voltage while it is connected to the PZT. I attached a plot of the actual voltage measured by the DVM vs the EPICS reading.
It shows a hysteresis.
Also the actual voltage drops by more than a half when the PZT is connected. The output impedance of the HV amp is 64k (according to the schematic). If I believe this number, the impedance of the PZT should also be 64k. The current flowing the PZT is 1.6mA at 200V EPICS reading.
The impedance of the PZT directly measured by the DVM is 1.5M ohm, which is significantly different from the value expected above. I will check the actual output impedance of the HV amp later.
The capacitance of the PZT measured by the DVM is 300nF. I don't know if I can believe the DVM's ability to measure C.

I noticed that when a high voltage is applied, the actual voltage across the PZT shows a decay.
The second plot shows the step response of the actual voltage.
The voltage coming to the PZT was T-ed and reduced by a factor of 30 using a high impedance voltage divider to be recorded by an ADC.
The PMCTRANSPD channel is temporarily used to monitor this signal.
After the voltage applied to the PZT was increased abruptly (to ~230V), the actual voltage starts to exponentially decrease.
When the HV was reduced to ~30V, the actual voltage goes up. This behavior explains the weird exponential motion of the PZT feedback signal when the PMC is locked.
The cause of the actual voltage drop is not understood yet.
From the above measurements, we can almost certainly conclude that the problem of the PMC is in the PZT, not in the HV amp nor the read back.


I'd believe the Fluke's measurement of capacitance. Here's some info from PK about the PZT:


Quote:

But the PMC ones were something like
0.750 in. thick x 0.287 in. thick. 2 microns per 200 V displacement,
resonant frequency greater than 65 kHz. Typical capacitance is around 0.66
uF.


If the PZT capacitance has dropped by a factor of two, that seems like a bad sign. I don't know what to expect for a resistance value of the PZT, but I wouldn't be surprised if it's non-Ohmic. The 64k is the series resistor after the PA85, not the modeled resistance of the PZT itself.
  700   Fri Jul 18 19:43:55 2008 YoichiDAQComputersPSL fast channels cannot be read by dataviewer
At this moment only the PSL fast channels have trouble.
Rob restarted fb40m, c1IOVME, but no effect.
  699   Fri Jul 18 19:41:09 2008 YoichiUpdatePSLPMC PZT investigation
I measured the HV coming to the PMC PZT by plugging it off from the PZT and hooking it up to a DVM.
The reading of DVM is pretty much consistent with the reading on EPICS. I got 287V on the DVM when the EPICS says 290V.

Then I used a T to monitor the same voltage while it is connected to the PZT. I attached a plot of the actual voltage measured by the DVM vs the EPICS reading.
It shows a hysteresis.
Also the actual voltage drops by more than a half when the PZT is connected. The output impedance of the HV amp is 64k (according to the schematic). If I believe this number, the impedance of the PZT should also be 64k. The current flowing the PZT is 1.6mA at 200V EPICS reading.
The impedance of the PZT directly measured by the DVM is 1.5M ohm, which is significantly different from the value expected above. I will check the actual output impedance of the HV amp later.
The capacitance of the PZT measured by the DVM is 300nF. I don't know if I can believe the DVM's ability to measure C.

I noticed that when a high voltage is applied, the actual voltage across the PZT shows a decay.
The second plot shows the step response of the actual voltage.
The voltage coming to the PZT was T-ed and reduced by a factor of 30 using a high impedance voltage divider to be recorded by an ADC.
The PMCTRANSPD channel is temporarily used to monitor this signal.
After the voltage applied to the PZT was increased abruptly (to ~230V), the actual voltage starts to exponentially decrease.
When the HV was reduced to ~30V, the actual voltage goes up. This behavior explains the weird exponential motion of the PZT feedback signal when the PMC is locked.
The cause of the actual voltage drop is not understood yet.
From the above measurements, we can almost certainly conclude that the problem of the PMC is in the PZT, not in the HV amp nor the read back.
Attachment 1: Hysteresis.png
Hysteresis.png
Attachment 2: StepResponse.png
StepResponse.png
  698   Fri Jul 18 19:30:20 2008 MashaUpdateAuxiliary lockingmoving from 40m
I will be working in the basement of Bridge probably starting next week; I moved the NPRO laser and some of the optics from my mach zehnder setup on the SP table to Bridge. Thanks for your help!
  697   Fri Jul 18 19:15:15 2008 JenneUpdateIOOChecking out the MC Servo Board

Quote:
[Jenne, Max]

I haven't checked to ensure that the MC still locks, because Yoichi is doing stuff on the PSL table, but I didn't change anything on the board, and hooked all the cables back where they were, so hopefully it's all okay.


I put the PMC back and the MC now locks.
  696   Fri Jul 18 17:12:35 2008 JenneUpdateIOOChecking out the MC Servo Board
[Jenne, Max]

One of the things that Rana thinks that might be causing my MC_F calibration to be off is that the MC Servo Board's filters don't match those on the schematics. Max and I pulled the MC servo board today to check resistor and capacitor values. Alberto needed the Mode Cleaner, so we put the board back before finishing checking values. We will probably pull the board again next week to finish checking the values.

I haven't checked to ensure that the MC still locks, because Yoichi is doing stuff on the PSL table, but I didn't change anything on the board, and hooked all the cables back where they were, so hopefully it's all okay.
  695   Fri Jul 18 17:06:20 2008 JenneUpdateComputersComputers down for most of the day, but back up now
[Sharon, Alex, Rob, Alberto, Jenne]

Sharon and I have been having trouble with the C1ASS computer the past couple of days. She has been corresponding with Alex, who has been rebooting the computers for us. At some point this afternoon, as a result of this work, or other stuff (I'm not totally sure which) about half of the computers' status lights on the MEDM screen were red. Alberto and Sharon spoke to Alex, who then fixed all of them except C1ASC. Alberto and I couldn't telnet into C1ASC to follow the restart procedures on the Wiki, so Rob helped us hook up a monitor and keyboard to the computer and restart it the old fashioned way.

It seems like C1ASC has some confusion as to what its IP address is, or some other computer is now using C1ASC's IP address.

As of now, all the computers are back up.
  694   Fri Jul 18 16:57:37 2008 JenneUpdateIOOCalibrated MC_F
I have calibrated MC_F. The conversion factor is 137.49 MHz/count.

The calibration data taken is attached, along with a calibrated power spectrum.

On the data plot, the x axis is volts from the C1:IOO-MC_FAST_MON channel, with the calibration between FAST_MON and MC_F = -788.18 volts/count.
The linear term of the fit line = -0.085MHz/volt. Error bars are +/- 1 in the last digit of what the spectrum analyzer gave me for frequency (+/- 0.01MHz).

The net conversion factor is then (-788.18)*(-0.085)*(2) = 137.49 MHz/count. The factor of 2 is because the light passes through the AOM twice.

On the power spectrum,
REF0 and REF1 = MC unlocked, HEPAs on, MC Refl gain = 22
REF2 and REF3 = MC locked, HEPAs on, MC Refl gain = 22
REF4 and REF5 = MC locked, HEPAs on, MC Refl gain = 19
REF6 and REF7 = MC locked, HEPAs off, MC Refl gain = 19
Attachment 1: MC_Fcalib.png
MC_Fcalib.png
Attachment 2: 20080717MC_F-MC_I.pdf
20080717MC_F-MC_I.pdf
  693   Fri Jul 18 12:24:15 2008 josephb, EricConfigurationCamerasChanged Lenses on GC750 at ETMX
We removed the giant TV zoom lens and replaced it with a much smaller fixed zoom lens. Currently it views the entire optic. We have another (also small) zoom lens which focuses much better on the spot itself. With how far back the camera is currently placed, neither of these fixed zoom lenses can touch or hit the view port or the chamber while still attached to camera and mount, even using all of the mount's motion range. So this should be less of a safety issue.

Ideally, we'd like to get some images of the full optic (including osems and so forth) with the X-arm locked, and then use the higher zoom lens while still locked, to get images we can use to calibrate the x and y length scales.
  692   Thu Jul 17 20:13:34 2008 YoichiUpdatePSLPMC alignment/mode matching effort
I'm working to improve the mode matching of PMC.
Because I noticed that the beam was hitting the aperture of the EOM for PMC, I moved the EOM a little bit to maximize the transmission.
This did not change the alignment to the reference cavity but changed the alignment of the PMC a lot.
So I adjusted it back.
The alignment of the PMC can be easily optimized but the Hermite 02 mode still remains. This means the mode matching is bad.
Moving the lenses by a small amount (a few mm) did not change the height of 02 mode.
I'm planning to move the lenses by a large amount tomorrow. But it will destroy the alignment to the PMC.
So I installed two irises in the beam path after the lenses to remember the alignment roughly.
Right now the PMC transmission is slightly worse than before because the lens positions are not good.
  691   Thu Jul 17 16:39:58 2008 Max JonesUpdateDAQMagnetometer Installed
Today I installed the magnetometer near the beam splitter chamber. It is located on the BSC chamber at head height on the inner part of the interferometer (meaning I had to crawl under the arms to install it). I don't think I disturbed anything during installation but I think that's it's probably prudent to tell everyone that I was back there just in case. I plan to run 3 BNC cables (one for each axis) from the magnetometer to the DAQ input either tonight or tomorrow. Suggestions are appreciated. - Max.
  690   Thu Jul 17 13:08:37 2008 JohnSummaryLSCHOM resonances in the arms
On Tuesday night we attempted to lock the full DRFPMI in the optical spring configuration with the +f2 sideband resonant in the SRC.
Despite having no problems locking on the +f2 in a DRM we couldn't lock the full IFO.

There was some discussion about whether a +f2 higher order mode resonance in the arms could cause this problem.

I calculated the positions of the first six higher order modes for the carrier and all sidebands (using Siegman p 762 (23) with a plus sign).
Plot attached. Colors indicate different frequency components, numbers are the mode index (m+n). Thick lines are fundamental modes of
the sidebands. Heights of HOM indicators have been scaled by 1/(m+n)^2.

It appears that the first order transverse mode of the +166 is indeed partially resonant. We might try to tweak the sideband frequencies a
little to see if this helps us. It would probably be prudent to measure the MC length first.

Numbers used:

L = 38.5750; %average of Alberto's recent measurements elog #556
Retm = 57.375;
f166 = 165.977195e6;
f33 = 33.195439e6;
Attachment 1: HOMresonances.png
HOMresonances.png
  689   Thu Jul 17 12:15:21 2008 EricUpdatePSLSwept PMC PZT voltage range
I unlocked the PMC and swept over C1:PSL-PMC_RAMP's full range a couple of times this morning.  The PMC should now be relocked and returned 
to normal.
  688   Thu Jul 17 08:30:15 2008 steveUpdatePSLPMC relocked manually
The PMC pzt HV and the servo gain adj. are railing at max this morning

Why is it on the decreasing side of FSS_RTTEMP slope?
Attachment 1: pmc4d.jpg
pmc4d.jpg
  687   Thu Jul 17 00:59:18 2008 JenneSummaryGeneralFunny signal coming out of VCO
While working on calibrating the MC_F signal, Rana and I noticed a funny signal coming out of the VCO. We expect the output to be a nice sine wave at about 80MHz. What we see is the 80MHz signal plus higher harmonics. The reason behind the craziness is to be determined. For now, here's what the signal looks like, in both time and frequency domains.

The first plot is a regular screen capture of a 'scope. The second is the output of the SR spectrum analyzer, as seen on a 'scope screen. The leftmost tall peak is the 80MHz peak, and the others are the harmonics.
Attachment 1: VCOout_time.PNG
VCOout_time.PNG
Attachment 2: VCOout_freq.PNG
VCOout_freq.PNG
  686   Wed Jul 16 22:29:05 2008 MashaUpdateAuxiliary lockinglong measurement

Quote:
I'm taking a measurement on the SR785 spectrum analyzer at low frequencies, so I'm going to leave it by the symmetric port table for a while. Please don't move it!


all done thanks.
  685   Wed Jul 16 17:51:58 2008 MashaUpdateAuxiliary lockinglong measurement
I'm taking a measurement on the SR785 spectrum analyzer at low frequencies, so I'm going to leave it by the symmetric port table for a while. Please don't move it!
  684   Wed Jul 16 17:36:51 2008 JohnDAQPSLFSS input offset
I changed the nominal FSS input offset to 0 from 0.3. Tolerance remains unchanged at +/-0.05.
ELOG V3.1.3-